# Exploring Human-in-the-Loop Test-Time Adaptation by Synergizing Active Learning and Model Selection

### v2

@@ Appendix B Additional Experimental Analysis
Computation Complexity of Competing Active TTA Methods: We find that existing stream-based active learning (VeSSAL (Kossen et al., 2022)) and active TTA (SimATTA (Gui et al., 2024)) are inherently more computationally expensive than HILTTA due to the following reasons. First, VeSSAL adopts gradient embedding, i.e. $g(x_{i})=\partial l(f(x_{i};\theta),y_{i})/\partial\theta_{L}$, where $\theta_{L}\in\mathbb{R}^{K\times D}$ refers to the classifier weights. This results in a very high dimensional gradient embedding $g(x_{i})\in\mathbb{R}^{D\cdot K}$. The probability of labeling a sample $p_{i}$ involves calculating the inverse of covariance matrix, i.e. $p_{i}\propto g(x_{i})^{\top}\hat{\Sigma}^{-1}g(x_{i}),\;s.t.\hat{\Sigma}=\sum_{x_{i}\in\mathcal{B}}g(x_{i})g(x_{i})^{\top}\in\mathbb{R}^{DK\times DK}$. Hence, VeSSAL can hardly deal with classification tasks with a large number of classes, e.g. ImageNet, where $K$ is large and calculating a large matrix inverse is inherently time-consuming. Second, SimATTA employs a K-Means clustering algorithm for the selection of samples to annotate. Since K-Means is an iterative approach the computation cost of SimATTA is relatively high as well. In contrast, our proposed sample selection is built upon K-Center clustering which sequentially selects the sample with the highest distance to cluster centers (greedy algorithm), the computation cost is significantly lower. An empirical study in Tab. 10 reveals that our method is about 3 times faster than VeSSAL and SimATTA in adaptation time.
@@
Random Batchsizes: We recognize that in certain scenarios, test data may arrive inconsistently and intermittently. Results from Tab. 11 demonstrate that our proposed HILTTA framework is adaptable to such varying batchsize. At each iteration, we vary the batchsize randomly among the choices [50, 100, 150, 200]. This versatility leads to even greater improvements compared to fixed batch size. This enhanced performance stems from the fact that the optimal hyper-parameter values vary with different batch sizes, thus benefiting significantly from our adaptive model selection approach.
@@
Ablation Study with RMT: We have conducted wider experiments to analyze the ablation study with RMT (Döbler et al., 2023) as the off-the-shelf TTA method in Tab. 12 exhibits a consistent trend with the findings presented in the manuscript, providing additional support for our analysis and demonstrating the effectiveness of our proposed approach.

### v3

@@ B.6 Comparing Adaptation Time & Annotation Time
B.7 Computation Complexity of Competing Active TTA Methods
We find that existing stream-based active learning (VeSSAL (Kossen et al., 2022)) and active TTA (SimATTA (Gui et al., 2024)) are inherently more computationally expensive than HILTTA due to the following reasons. First, VeSSAL adopts gradient embedding, i.e. $g(x_{i})=\partial l(f(x_{i};\theta),y_{i})/\partial\theta_{L}$, where $\theta_{L}\in\mathbb{R}^{K\times D}$ refers to the classifier weights. This results in a very high dimensional gradient embedding $g(x_{i})\in\mathbb{R}^{D\cdot K}$. The probability of labeling a sample $p_{i}$ involves calculating the inverse of covariance matrix, i.e. $p_{i}\propto g(x_{i})^{\top}\hat{\Sigma}^{-1}g(x_{i}),\;s.t.\hat{\Sigma}=\sum_{x_{i}\in\mathcal{B}}g(x_{i})g(x_{i})^{\top}\in\mathbb{R}^{DK\times DK}$. Hence, VeSSAL can hardly deal with classification tasks with a large number of classes, e.g. ImageNet, where $K$ is large and calculating a large matrix inverse is inherently time-consuming. Second, SimATTA employs a K-Means clustering algorithm for the selection of samples to annotate. Since K-Means is an iterative approach the computation cost of SimATTA is relatively high as well. In contrast, our proposed sample selection is built upon K-Center clustering which sequentially selects the sample with the highest distance to cluster centers (greedy algorithm), the computation cost is significantly lower. An empirical study in Tab. 13 reveals that our method is about 3 times faster than VeSSAL and SimATTA in adaptation time.
@@ B.7 Computation Complexity of Competing Active TTA Methods
B.8 Studies on Random Batchsizes
We recognize that in certain scenarios, test data may arrive inconsistently and intermittently. Results from Tab. 14 demonstrate that our proposed HILTTA framework is adaptable to such varying batchsize. At each iteration, we vary the batchsize randomly among the choices [50, 100, 150, 200]. This versatility leads to even greater improvements compared to fixed batch size. This enhanced performance stems from the fact that the optimal hyper-parameter values vary with different batch sizes, thus benefiting significantly from our adaptive model selection approach.
@@ B.8 Studies on Random Batchsizes
B.9 Further Ablation Study with RMT
We have conducted wider experiments to analyze the ablation study with RMT (Döbler et al., 2023) as the off-the-shelf TTA method in Tab. 15 exhibits a consistent trend with the findings presented in the manuscript, providing additional support for our analysis and demonstrating the effectiveness of our proposed approach.